# SurvMamba: State Space Model with Multi-grained Multi-modal Interaction for Survival Prediction

### v1

@@ Abstract.
Multi-modal learning, Survival Prediction, Multi-grained, Mamba
ccs: Applied computing Life and medical sciences
1. Introduction
@@
Survival prediction evaluates patients’ mortality risks, thereby enhancing the clinical decision-making process related to diagnosis and treatment planning (Tran et al., 2021). For cancer patients, pathological images and genomic profiles provide critical and interconnected information for patient stratification and survival analysis (Shao et al., 2019). For example, pathological images detail the tumor microenvironment, capturing the diversity of cancer cells and immune interactions (Gui et al., 2023). At the same time, genomic profiles provide critical insights into cancer cell states and immune system factors that affect dynamic prognostic outcomes (Milanez-Almeida et al., 2020; Tong et al., 2023). Therefore, integrating pathological images and genomic data through multi-modal learning holds considerable promise to enhance the precision of cancer survival predictions.
@@
Extensive efforts have been made in multi-modal survival analysis with histological whole slide images (WSIs) and transcriptomic data (Jaume et al., 2023; Xu and Chen, 2023; Subramanian et al., 2021). Among them, Multiple Instance Learning (MIL) (Jaume et al., 2023; Xu and Chen, 2023; Qiu et al., 2023) has emerged as an effective method to process the high-dimensional WSIs and transcriptomic data. In these MIL-based methods, each WSI is represented as a “bag” with numerous patches as instances, while transcriptomic data is organized into a “bag” with functional genomes (i.e., genomic groups) as instances. Subsequently, fusion of histological and genomic features (Li et al., 2022; Chen et al., 2022b) is conducted to predict survival outcomes.
@@
Despite the complex, high-dimensional nature of WSIs and transcriptomic data, they exhibit significant inherent hierarchical structures. These structures stem from their fundamental biological functions and the pathology associated with diseases, as illustrated in Figure 1a. However, existing methods do not fully leverage these hierarchical structures in both WSIs and transcriptomic data (Chen et al., 2022a; Kanehisa and Goto, 2000), which may lead to the following issues. The first issue regards to insufficient global representation. There are multi-level prognostic insights reflected in the hierarchical information. For example, fine-grained patch-level pathological images detail cell densities (Zhang et al., 2024a), while coarse-grained region-level images reveal tissue features and tumor-immune interactions (Fu et al., 2020). Fine-grained function-level transcriptomics reveals the specific functionalities of gene sets (Ietswaart et al., 2021), while coarse-grained process-level data focus on identifying those fundamental macroscopic biological processes (Zhang et al., 2021). Therefore, important global features for survival prediction could potentially be lost if a method focuses on fine-grained information only. Furthermore, there is limited cross-modality communications. Both WSIs and transcriptomic data contain a multitude of cross-modality communications at different hierarchical levels. For instance, adjacent patches/regions within a WSI tend to exhibit higher correlations (Ke et al., 2023), and functions/process sharing similar or related biological mechanisms demonstrate closer relationships (Smug et al., 2023), which cannot be easily revealed if cross-modality communications are only established at fine-grained features level, leading to inadequate intra-modal representations and and subsequently deficient inter-modal integration.
@@
Although the integration of hierarchical structures into multi-modal survival prediction frameworks holds promise for yielding more comprehensive prognostic insights, it poses significant computational challenges, especially when implemented using attention mechanisms, given their high computational complexity. On the other hand, the Selective Structured State Space Model, known as Mamba (Gu and Dao, 2023), has demonstrated remarkable efficiency in long sequence modeling. Mamba effectively captures long-range dependencies and improves training and inference efficiency through a selection mechanism and a hardware-aware algorithm, hence providing an alternative to deal with high-dimensional data.
(a) Hierarchical representation of WSI and transcriptomics. Left: WSI reveals tissue organization at region-level and provides detailed cellular insights at patch-level. Right: Genes can be categorized according to their genomic function, which can be further subdivided based on biological process.
(b) Left: Attention-based methods focus solely on fine-grained information, resulting in high computational costs and could potentially miss crucial global information. Right: Our State Space Model (SSM)-based SurvMamba captures more comprehensive multi-modal information with efficient computation.
Figure 1. (a) Hierarchical structure of WSI and transcriptomic data. (b) Comparison of previous methods with SurvMamba.
@@
In this paper, we propose a Mamba-based survival prediction method with multi-grained multi-modal interaction, termed SurvMamba, which extracts multi-modal multi-grained information from hierarchical structure and facilitates efficient intra-modal and inter-modal interactions. Specifically, we design a novel Hierarchical Interaction Mamba (HIM) to efficiently capture the hierarchical characteristics of WSIs and transcriptomic data. HIM extracts coarse-grained features through the aggregation of fine-grained instances and enables efficient bidirectional interactions for multi-grained instances. In this way, SurvMamba can extract enhanced local information from fine-grained instances and global information from coarse-grained instances, resulting in more comprehensive intra-modal information to predict patient survival outcomes. Furthermore, we propose an Interaction Fusion Mamba (IFM) to facilitate interactions between histological and genomic features across different granularities, thereby providing refined intra-modal representations at both fine and coarse levels. Finally, these multi-grained features are adaptively integrated to formulate the final survival prediction. The main contributions of this paper can be summarized as follows:
@@
2. Related Work
2.1. Multi-modal Survival Prediction
@@
Survival outcome prediction, also known as time-to-event analysis (David and Mitchel, 2012), concentrates on probabilistic assessment of experiencing a specified event such as mortality in the clinical setting before a time under both uncensored and right-censored data. Right-censored data represents cases wherein the event of interest remains unobserved throughout the duration of the study. In the current state-of-the-art methods, estimating cancer patient survival heavily depends on physicians’ assessment of histology and/or interpretation of genomic sequencing report (Chen et al., 2021b). Consequently, there is a growing interest in multi-modal learning methods that integrate histopathology and genomic data for survival prediction (Chen et al., 2020, 2021a, 2021b; Subramanian et al., 2021; Chen et al., 2022b; Jaume et al., 2023; Xu and Chen, 2023). Chen et al. (Chen et al., 2021b) proposed a Multi-modal Co-Attention Transformer framework that identifies informative instances from pathological images using genomic features as queries. Qiu et al. (Qiu et al., 2023) proposed PONET, a novel pathology-genomic deep model informed by biological pathway that integrates pathological images and genomic data to improve survival prediction. Jaume et al. (Jaume et al., 2023) modeled interactions between biological pathway and histology patch tokens using a memory-efficient multi-modal Transformer for survival analysis. Zhang et al. (Zhang et al., 2024b) proposed a new framework named Prototypical Information Bottlenecking and Disentangling (PIBD), including Prototypical Information Bottleneck module for intra-modal redundancy and Prototypical Information Disentanglement module for inter-modal redundancy. Existing methods mainly regard multi-modal survival prediction as a fine-grained recognition task, focusing on histological patches and genomic functions to model fine-grained cross-modal relationships with local information. Neglecting coarser-grained information may result in overlooking some essential global features vital for accurately predicting patient survival. WSIs and genomic data exhibit an intrinsic hierarchical structure, where different granularity levels yield unique and critical insights into patient prognosis.
2.2. State Space Models
@@
Recently, the State Space Models (SSMs) have shown significant effectiveness of state space transformation in capturing the dynamics and dependencies of language sequences. The structured state-space sequence model (S4) introduced in (Gu et al., 2021) is specifically designed to model long-range dependencies, offering the advantage of linear complexity. Various models including S5 (Smith et al., 2022), H3 (Fu et al., 2022) and GSS (Mehta et al., 2022) have been developed base on S4, and Mamba (Gu and Dao, 2023) distinguishes itself by introducing a data-dependent SSM layer and a selection mechanism using parallel scan (S6). Compared to Transformers with quadratic-complexity attention, Mamba excels at processing long sequences with linear complexity. Existing models based on MIL for multi-modal survival prediction struggle with enabling effective and efficient interactions among vast numbers of instances (Ilse et al., 2018; Li et al., 2021; Shao et al., 2021; Li et al., 2024), burdened by significant computational requirements. The introduction of Mamba addresses these challenges by incorporating input-adaptive and global information modeling techniques that emulate self-attention functionalities while retaining linear complexity. This breakthrough diminishes the computational overhead and provides a potential multi-modal learning framework for survival prediction.
3. Methodology
@@
Figure 2. Overview of SurvMamba architecture. WSIs and transcriptomics are illustrated in a three-layer structure, comprising WSI/Region/Patch for WSIs and Gene/Process/Function for transcriptomics, respectively. Initially, Pathology and Genomics Encoders extract fine-grained features in groups. Then, the Hierarchical Interaction Mamba (HIM) module enhances intra-modal feature interactions across various granularities to improve unimodal representation. Meanwhile, the Interaction Fusion Mamba (IFM) module enables inter-modal integration across different levels for more comprehensive multi-modal representations. Finally, multi-grained multi-modal features are adaptively integrated to predict survival risk scores.
3.1. Preliminaries
@@
(1) $\begin{split}h^{\prime}(t)&=\mathbf{A}h(t)+\mathbf{B}x(t)\\ y(t)&=\mathbf{C}h(t)\end{split}$
@@
(2) $\begin{split}\mathbf{\bar{A}}&=\exp(\mathbf{\Delta}\mathbf{A})\\ \mathbf{\bar{B}}&=(\mathbf{\Delta}\mathbf{A})^{-1}(\exp(\mathbf{\Delta}\mathbf{A})-I)\cdot\mathbf{\Delta}\mathbf{B}.\end{split}$
@@
(3) $\begin{split}h_{t}&=\bar{\mathbf{A}}h_{t-1}+\bar{\mathbf{B}}x_{t},\\ y_{t}&=\mathbf{C}h_{t}.\end{split}$
@@
(4) $\begin{split}\mathbf{\bar{K}}&=(\mathbf{C}\mathbf{\bar{B}},\mathbf{C}\bar{\mathbf{AB}},\ldots,\mathbf{C}\bar{\mathbf{A}}^{M-1}\mathbf{\bar{B}}),\\ \mathbf{y}&=\mathbf{x}*\mathbf{\bar{K}}.\end{split}$
3.2. Overview and Problem Formulation
@@
To capture more comprehensive multi-modal representations, we exploit the inherent hierarchical structure of WSI and transcriptomic data to extract multi-grained information, as will be described in Sec 3.3, subsequently enabling efficient interaction and fusion among them through Mamba-based modules. Following (Jaume et al., 2023; Xu and Chen, 2023; Qiu et al., 2023), we adopt a MIL framework to learn pathological and genomic representations. We first formulate each WSI and transcriptomics as a “bag”, extracting features of fine-grained instances (Patch or Function) from the Pathology Encoder and the Genomics Encoder in groups. Then, with bidirectional Mamba in dual-level MIL framework, the Hierarchical Interaction Mamba (HIM) module aggregates fine-grained instances into coarse-grained instances (Region or Process) and enables efficient intra-modal interactions at different granularities. The details of this module will be described in Sec. 3.4. Further, the Interaction Fusion Mamba (IFM) module facilitates multi-grained interactions between histological and genomic features to derive fine-grained and coarse-grained inter-modal representations, as will be described in Sec. 3.5. Finally, multi-grained multi-modal features will be adaptively fused to predict a hazard function and get the survival risk scores, as will be described in Sec. 3.6.
3.3. Hierarchical Multi-modal Representations
@@
Given a set of transcriptomics measurements, it can be mapped into $J$ different biological processes $S$. The $j$-th process $S_{j}$ contains $K_{j}$ genomic functions $F$, where $G=\{S_{1},S_{2},\ldots,S_{J}\}$, and $S_{j}=\{F_{j1},F_{j2},\ldots,F_{jK_{j}}\},1\leq j\leq J$. The number of genomic functions contained within each process varies. We encode genomic functions with Genomics Encoder $f_{G}(\cdot)$ which contains multilayer perceptrons (MLPs) with learnable weights, resulting in function-level token sequence $T^{G}_{j}$ in $S_{j}$, $T^{G}_{j}=\{f_{G}(F_{j1}),f_{G}(F_{j2}),\ldots,f_{G}(F_{jK_{j}})\}$.
3.4. Hierarchical Interaction Mamba
@@
Structural WSI and transcriptomic data exhibit certain dependencies within local or global features. To capture these dependencies, we develop the HIM module. In inspired by (Zhu et al., 2024), this module is designed to learn intra-model features by integrating a bidirectional Mamba (Bi-Mamba) mechanism within a dual-level MIL framework. A comprehensive overview of the Bi-Mamba is provided in Algorithm 1. Our approach aims to model correlations across features of varying granularities effectively. For the first level MIL, the HIM module regards fine-grained (i.e., patch- and function-level) features $T^{I}_{i}$ and $T^{G}_{j}$ as instances, using Bi-Mamba with shared parameters between groups (i.e., region or process) to model fine-grained histological and genomic long sequences, resulting in enhanced fine-grained features ${T^{I}_{m}}^{\prime}$ and ${T^{G}_{j}}^{\prime}$:
(5) $\begin{split}{T^{I}_{m}}^{\prime}&=\textbf{Bi-Mamba}(T^{I}_{m})\\ {T^{G}_{j}}^{\prime}&=\textbf{Bi-Mamba}(T^{G}_{j}).\end{split}$
@@
Subsequently, with the second level MIL, features ${T^{I}_{m}}^{\prime}$ and ${T^{G}_{j}}^{\prime}$ are pooling into coarse-grained (i.e., region- and process-level) features, and Bi-Mamba is then re-applied to facilitate interactions among them to get improved coarse-grained features $T^{I}$ and $T^{G}$. Through this dual-level deep interaction mechanism for each modality, the HIM ensures a thorough and efficient integration of local fine-grained and global coarse-grained features, obtaining enhanced and comprehensive intra-modal histological and genomic features.
@@
(6) $\begin{split}T^{I}&=\textbf{Bi-Mamba}([\text{Pool}({T^{I}_{1}}^{\prime}),\text{Pool}({T^{I}_{2}}^{\prime}),\ldots,\text{Pool}({T^{I}_{M}}^{\prime})])\\ T^{G}&=\textbf{Bi-Mamba}([\text{Pool}({T^{G}_{1}}^{\prime}),\text{Pool}({T^{G}_{2}}^{\prime}),\ldots,\text{Pool}({T^{G}_{J}}^{\prime})]).\end{split}$
Algorithm 1 Bi-Mamba
1:Input: token sequence $T_{l-1}:(B,M,D)$
2:Output: token sequence $T_{l}:(B,M,D)$
3:$T^{\prime}_{l-1}:(B,M,D)\leftarrow\text{Norm}(T_{l-1})$
4:$x:(B,M,E)\leftarrow\text{Linear}^{x}(T^{\prime}_{l-1})$
5:$z:(B,M,E)\leftarrow\text{Linear}^{z}(T^{\prime}_{l-1})$
6:for o in {forward, backward} do
7:     $x^{\prime}_{o}:(B,M,E)\leftarrow\text{SiLU}(\text{Conv1d}_{o}(x))$
8:     $B_{o}:(B,M,N)\leftarrow\text{Linear}^{B}_{o}(x^{\prime}_{o})$
9:     $C_{o}:(B,M,N)\leftarrow\text{Linear}^{C}_{o}(x^{\prime}_{o})$
10:     $\Delta_{o}:(B,M,E)\leftarrow\log(1+\exp(\text{Linear}^{\Delta}_{o}(x^{\prime}_{o})+\text{Parameter}^{\Delta}_{o}))$
11:     $\overline{A_{o}}:(B,M,E,N)\leftarrow\Delta_{o}\otimes\text{Parameter}^{A}_{o}$
12:     $\overline{B_{o}}:(B,M,E,N)\leftarrow\Delta_{o}\otimes B_{o}$
13:     $Y_{o}:(B,M,E)\leftarrow\text{SSM}(\overline{A_{o}},\overline{B_{o}},C_{o})(x^{\prime}_{o})$
14:end for
15:$Y^{\prime}_{\text{forward}}:(B,M,E)\leftarrow Y_{\text{forward}}\odot\text{SiLU}(z)$
16:$Y^{\prime}_{\text{backward}}:(B,M,E)\leftarrow Y_{\text{backward}}\odot\text{SiLU}(z)$
17:$T_{l}:(B,M,D)\leftarrow\text{Linear}^{T}(Y^{\prime}_{\text{forward}}+Y^{\prime}_{\text{backward}})+T_{l-1}$
18:return $T_{l}$
3.5. Interaction Fusion Mamba

### v2

@@ Abstract
1 Introduction
@@
Survival prediction evaluates patients’ mortality risks, thereby enhancing the clinical decision-making process related to diagnosis and treatment planning Tran et al. (2021). For cancer patients, pathological images and genomic profiles provide critical and interconnected information for patient stratification and survival analysis Shao et al. (2019). For example, pathological images detail the tumor microenvironment, capturing the diversity of cancer cells and immune interactions Gui et al. (2023). At the same time, genomic profiles provide critical insights into cancer cell states and immune system factors that affect dynamic prognostic outcomes Milanez-Almeida et al. (2020); Tong et al. (2023). Therefore, integrating pathological images and genomic data through multi-modal learning holds considerable promise to enhance the precision of cancer survival predictions.
@@
Extensive efforts have been made in multi-modal survival analysis with histological whole slide images (WSIs) and transcriptomic data Jaume et al. (2023); Xu and Chen (2023); Subramanian et al. (2021). Among them, Multiple Instance Learning (MIL) Jaume et al. (2023); Xu and Chen (2023); Qiu et al. (2023) has emerged as an effective method to process the high-dimensional WSIs and transcriptomic data. In these MIL-based methods, each WSI is represented as a “bag” with numerous patches as instances, while transcriptomic data is organized into a “bag” with functional genomes (i.e., genomic groups) as instances. Subsequently, fusion of histological and genomic features Li et al. (2022); Chen et al. (2022b) is conducted to predict survival outcomes.
@@
Despite the complex, high-dimensional nature of WSIs and transcriptomic data, they exhibit significant inherent hierarchical structures. These structures stem from their fundamental biological functions and the pathology associated with diseases, as illustrated in Figure 1. WSI reveals tissue organization at region-level and provides detailed cellular insights at patch-level. Genes can be categorized according to their genomic function, which can be further subdivided based on biological process. However, existing methods do not fully leverage these hierarchical structures in both WSIs and transcriptomic data Chen et al. (2022a); Kanehisa and Goto (2000), which may lead to the following issues. The first issue regards to insufficient global representation. There are multi-level prognostic insights reflected in the hierarchical information. For example, fine-grained patch-level pathological images detail cell densities Zhang et al. (2024a), while coarse-grained region-level images reveal tissue features and tumor-immune interactions Fu et al. (2020). Fine-grained function-level transcriptomics reveals the specific functionalities of gene sets Ietswaart et al. (2021), while coarse-grained process-level data focus on identifying those fundamental macroscopic biological processes Zhang et al. (2021). Therefore, important global features for survival prediction could potentially be lost if a method focuses on fine-grained information only. Furthermore, there is limited cross-modality communications. Both WSIs and transcriptomic data contain a multitude of cross-modality communications at different hierarchical levels, including patch-to-function and region-to-process relationships Ke et al. (2023); Smug et al. (2023), which cannot be easily revealed if cross-modality communications are only established at fine-grained features level, leading to deficient inter-modal integration.
@@
Although the integration of hierarchical structures into multi-modal survival prediction frameworks holds promise for yielding more comprehensive prognostic insights, it poses significant computational challenges, especially when implemented using attention mechanisms, given their high computational complexity. On the other hand, the Selective Structured State Space Model (SSM), known as Mamba Gu and Dao (2023), has demonstrated remarkable efficiency in long sequence modeling. Mamba effectively captures long-range dependencies and improves training and inference efficiency through a selection mechanism and a hardware-aware algorithm, hence providing an alternative to deal with high-dimensional data.
Figure 1: Hierarchical structure of WSI and transcriptomic data.
@@
In this paper, we propose a Mamba-based survival prediction method with multi-grained multi-modal interaction, termed SurvMamba, which extracts multi-modal multi-grained information from hierarchical structure and facilitates efficient intra-modal and inter-modal interactions. Specifically, we design a novel Hierarchical Interaction Mamba (HIM) to efficiently capture the hierarchical characteristics of WSIs and transcriptomic data. HIM extracts coarse-grained features through the aggregation of fine-grained instances and enables efficient bidirectional interactions for multi-grained instances. In this way, SurvMamba can extract enhanced local information from numerous fine-grained instances and global information from coarse-grained instances, resulting in more comprehensive intra-modal information to predict patient survival outcomes. Furthermore, we propose an Interaction Fusion Mamba (IFM) to facilitate interactions between histological and genomic features across different granularities, thereby providing refined intra-modal representations at both fine and coarse levels. Finally, these multi-grained features are adaptively integrated to formulate the final survival prediction. The main contributions of this paper can be summarized as follows:
@@
2 Related Work
2.1 Multi-modal Survival Prediction
@@
Survival outcome prediction, also known as time-to-event analysis David and Mitchel (2012), concentrates on probabilistic assessment of experiencing a specified event such as mortality in the clinical setting before a time under both uncensored and right-censored data. Right-censored data represents cases wherein the event of interest remains unobserved throughout the duration of the study. In the current state-of-the-art methods, estimating cancer patient survival heavily depends on physicians’ assessment of histology and/or interpretation of genomic sequencing report Chen et al. (2021b). Consequently, there is a growing interest in multi-modal learning methods that integrate histopathology and genomic data for survival prediction Chen et al. (2020, 2021a, 2021b); Subramanian et al. (2021); Chen et al. (2022b); Jaume et al. (2023); Xu and Chen (2023). Chen et al. Chen et al. (2021b) proposed a Multi-modal Co-Attention Transformer framework that identifies informative instances from pathological images using genomic features as queries. Qiu et al. Qiu et al. (2023) proposed PONET, a novel pathology-genomic deep model informed by biological pathway that integrates pathological images and genomic data to improve survival prediction. Jaume et al. Jaume et al. (2023) modeled interactions between biological pathway and histology patch tokens using a memory-efficient multi-modal Transformer for survival analysis. Zhang et al. Zhang et al. (2024b) proposed a new framework named Prototypical Information Bottlenecking and Disentangling (PIBD), including Prototypical Information Bottleneck module for intra-modal redundancy and Prototypical Information Disentanglement module for inter-modal redundancy. Existing methods mainly regard multi-modal survival prediction as a fine-grained recognition task, focusing on histological patches and genomic functions to model fine-grained cross-modal relationships with local information. Neglecting coarser-grained information may result in overlooking some essential global features vital for accurately predicting patient survival. WSIs and genomic data exhibit an intrinsic hierarchical structure, where different granularity levels yield unique and critical insights into patient prognosis.
2.2 State Space Models
@@
Recently, the State Space Models (SSMs) have shown significant effectiveness of state space transformation in capturing the dynamics and dependencies of language sequences. The structured state-space sequence model (S4) introduced in Gu et al. (2021) is specifically designed to model long-range dependencies, offering the advantage of linear complexity. Various models including S5 Smith et al. (2022), H3 Fu et al. (2022) and GSS Mehta et al. (2022) have been developed base on S4, and Mamba Gu and Dao (2023) distinguishes itself by introducing a data-dependent SSM layer and a selection mechanism using parallel scan (S6). Compared to Transformers with quadratic-complexity attention, Mamba excels at processing long sequences with linear complexity. Existing models based on MIL for multi-modal survival prediction struggle with enabling effective and efficient interactions among vast numbers of instances Ilse et al. (2018); Li et al. (2021); Shao et al. (2021); Li et al. (2024), burdened by significant computational requirements. The introduction of Mamba addresses these challenges by incorporating input-adaptive and global information modeling techniques that emulate self-attention functionalities while retaining linear complexity. This breakthrough diminishes the computational overhead and provides a potential multi-modal learning framework for survival prediction.
3 Methodology
@@
3.1 Preliminaries
@@
$\begin{split}h^{\prime}(t)&=\mathbf{A}h(t)+\mathbf{B}x(t)\\ y(t)&=\mathbf{C}h(t)\end{split}$ (1)
@@
$\begin{split}\mathbf{\bar{A}}&=\exp(\mathbf{\Delta}\mathbf{A})\\ \mathbf{\bar{B}}&=(\mathbf{\Delta}\mathbf{A})^{-1}(\exp(\mathbf{\Delta}\mathbf{A})-I)\cdot\mathbf{\Delta}\mathbf{B}.\end{split}$ (2)
@@
$\begin{split}h_{t}&=\bar{\mathbf{A}}h_{t-1}+\bar{\mathbf{B}}x_{t},\\ y_{t}&=\mathbf{C}h_{t}.\end{split}$ (3)
@@
$\begin{split}\mathbf{\bar{K}}&=(\mathbf{C}\mathbf{\bar{B}},\mathbf{C}\bar{\mathbf{AB}},\ldots,\mathbf{C}\bar{\mathbf{A}}^{M-1}\mathbf{\bar{B}}),\\ \mathbf{y}&=\mathbf{x}*\mathbf{\bar{K}}.\end{split}$ (4)
Figure 2: Overview of SurvMamba architecture. WSIs and transcriptomics are illustrated in a three-layer structure, comprising WSI/Region/Patch for WSIs and Gene/Process/Function for transcriptomics, respectively.
3.2 Overview and Problem Formulation
@@
To capture more comprehensive multi-modal representations, we exploit the inherent hierarchical structure of WSI and transcriptomic data to extract multi-grained information, subsequently enabling efficient interaction and fusion among them through Mamba-based modules. We adopt the MIL framework to learn pathological and genomic representations, formulating each WSI and transcriptomics as a “bag”. We extract features of fine-grained instances (Patch or Function) from the Pathology Encoder and the Genomics Encoder in groups. Then, with bidirectional Mamba in dual-level MIL framework, the Hierarchical Interaction Mamba (HIM) module aggregates fine-grained instances into coarse-grained instances (Region or Process) and enables efficient intra-modal interactions at different granularities. Further, the Interaction Fusion Mamba (IFM) module facilitates multi-grained interactions between histological and genomic features to derive fine-grained and coarse-grained inter-modal representations. Finally, multi-grained multi-modal features will be adaptively fused to predict a hazard function and get the survival risk scores.
3.3 Hierarchical Multi-modal Representations
@@
Given a set of transcriptomics measurements $G$, it can be mapped into $J$ different biological processes $S$. The $j$-th process $S_{j}$ contains $K_{j}$ genomic functions $F$, where $G=\{S_{1},S_{2},\ldots,S_{J}\}$, and $S_{j}=\{F_{j1},F_{j2},\ldots,F_{jK_{j}}\},1\leq j\leq J$. The number of genomic functions contained within each process varies. We encode genomic functions with Genomics Encoder $f_{G}(\cdot)$ which contains multilayer perceptrons (MLPs) with learnable weights, resulting in function-level token sequence $T^{G}_{j}$ in $S_{j}$, $T^{G}_{j}=\{f_{G}(F_{j1}),f_{G}(F_{j2}),\ldots,f_{G}(F_{jK_{j}})\}$.
3.4 Hierarchical Interaction Mamba
@@
Structural WSI and transcriptomic data exhibit certain dependencies within local or global features. To capture these dependencies, we develop the HIM module. Inspired by Zhu et al. (2024), this module is designed to learn intra-model features by integrating a bidirectional Mamba (Bi-Mamba) mechanism within a dual-level MIL framework. The specific structure of Bi-Mamba is shown in the top-right corner of Figure 2. Our approach aims to model correlations across features of varying granularities effectively. For the first level MIL, the HIM module regards fine-grained (i.e., patch- and function-level) features $T^{I}_{i}$ and $T^{G}_{j}$ as instances, using Bi-Mamba with shared parameters between groups (i.e., region or process) to model fine-grained histological and genomic long sequences, resulting in enhanced fine-grained features ${T^{I}_{m}}^{\prime}$ and ${T^{G}_{j}}^{\prime}$:
$\begin{split}{T^{I}_{m}}^{\prime}&=\textbf{Bi-Mamba}(T^{I}_{m})\\ {T^{G}_{j}}^{\prime}&=\textbf{Bi-Mamba}(T^{G}_{j}).\end{split}$ (5)
@@
Subsequently, with the second level MIL, features ${T^{I}_{m}}^{\prime}$ and ${T^{G}_{j}}^{\prime}$ are aggregated into coarse-grained (i.e., region- and process-level) features by average pooling, and Bi-Mamba is then re-applied to facilitate interactions among them to get improved coarse-grained features $T^{I}$ and $T^{G}$. Through this dual-level deep interaction mechanism for each modality, the HIM ensures a thorough and efficient integration of local fine-grained and global coarse-grained features, obtaining enhanced and comprehensive intra-modal histological and genomic features.
@@
$\begin{split}T^{I}&=\textbf{Bi-Mamba}([\text{Pool}({T^{I}_{1}}^{\prime}),\text{Pool}({T^{I}_{2}}^{\prime}),\ldots,\text{Pool}({T^{I}_{M}}^{\prime})])\\ T^{G}&=\textbf{Bi-Mamba}([\text{Pool}({T^{G}_{1}}^{\prime}),\text{Pool}({T^{G}_{2}}^{\prime}),\ldots,\text{Pool}({T^{G}_{J}}^{\prime})]).\end{split}$ (6)
3.5 Interaction Fusion Mamba